Cube root of 25x multiply by cube root of 9x
Answer by stanbon(75887)

Cube root of 25x multiply by cube root of 9x
--------------
= cbrt(25x*9x) = cbrt(225x^2)
------------
Cheers,
Stan H.
--------------